MICROWAVE DETECTION OF PROTONATED SO_{2} IN TWO ISOMERIC FORMS

\underline{V. LATTANZI}, M.C. MCCARTHY, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; SVEN THORWIRTH, Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, 53121 Bonn, Germany.

By means of Fabry-Pérot FT microwave spectroscopy, the rotational spectrum of protonated sulfur dioxide in two distinct isomeric forms, a cis and a trans geometry, is reported. Searches for both isomers were based on theoretical structures obtained at the CCSD(T)/cc-pwCVQZ level of theory corrected for zero-point vibrational effects at the CCSD(T)/ccpV(T+d)Z level. At a similarly high level of theory, the cis isomer is calculated to be the global minimum on the potential energy surface, but the trans isomer is predicted to lie only a few kcal/mol higher in energy. The identification of cis HOSO^{+} has been confirmed by detection of rotational lines of DOSO^{+} and HO^{34}SO^{+} at precisely the expected frequency shifts. Because sulfur dioxide is a well known interstellar and extragalactic species, and because it possesses a large proton affinity, HOSO^{+} is a excellent candidate for radioastronomical detection.